# SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense

### v2

@@ Appendix B MixUp
Appendix C Interval Neural Network Layers
@@
Interval Batch Normalization
@@
More concretely, given pre-activation interval bounds $[\underline{x},\overline{x}]$, we form the concatenated tensor:
@@
Interval Pooling Layers
@@
and the resulting output interval is $[\mathrm{MaxPool}_{\inf},\mathrm{MaxPool}_{\sup}]$.
@@
so that the final interval is $[\mathrm{AvgPool}_{\inf},\mathrm{AvgPool}_{\sup}]$.
@@
Appendix D Benchmark Datasets
Permuted MNIST introduces a different fixed pixel permutation for each task, applied to the original MNIST images. Rotated MNIST, on the other hand, creates tasks by rotating the digits by varying angles. These two datasets are widely used in the continual learning literature as synthetic benchmarks derived from MNIST (Goodfellow, Shlens, and Szegedy 2014; Liu and Liu 2022). Split CIFAR-100 (Zenke, Poole, and Ganguli 2017) is constructed by randomly partitioning the 100 classes of the CIFAR-100 dataset into 10 mutually exclusive groups, with each group forming a separate classification task of 10 classes. Similarly, Split miniImageNet is generated by selecting a subset of the original ImageNet dataset (Chaudhry et al. 2020) and dividing it into 20 distinct groups, each containing 5 unique classes. For both Split CIFAR-100 and Split miniImageNet, each class appears in only one group, ensuring no class overlap between tasks. This design allows each subset to be treated as an independent classification task, resulting in a sequence of tasks for continual learning. To match the setup of Split miniImageNet presented in (Ru et al. 2024), we split Split miniImageNet as described; however, we use only the first ten tasks. Therefore, we have 10 tasks, each with 5 classes.
Appendix E Discussion on Learnable Number of Parameters Used in SHIELD
@@
For the Permuted MNIST dataset, we reduce the number of learnable parameters to match the budget used in (Ru et al. 2024) by employing a hypernetwork composed of a two-layer MLP with hidden dimensions 100 and 50, and an embedding size of 96. The target network is a two-layer MLP with 50 neurons per layer. For the Rotated MNIST dataset, we adopt the same architecture but reduce the embedding size to 24 to further minimize the parameter count. We use $\varepsilon=\frac{25}{255}$ for both Permuted MNIST and Rotated MNIST during training.
For the Split CIFAR-100 dataset, again we align with the parameter budget from (Ru et al. 2024) by using an embedding size of 512 and a hypernetwork consisting of a two-layer MLP with 200 and 50 hidden units, respectively. Additionally, we simplify the target AlexNet architecture by removing two fully connected layers, applying a $\times 4$ 4 average pooling operation after the final convolutional layer in place of the max pooling, and reducing the hidden dimension of the last linear layer to 100. For Split CIFAR-100, we use $\varepsilon=\frac{2}{255}$ during training.
The total number of learnable parameters used in each configuration is reported in Table 2. Furthermore, we adopt the same FGSM, PGD, and AutoAttack settings as described in the main paper.
Table 2: Comparison of learnable parameters (in millions) between our method and the baseline methods from (Ru et al. 2024), which we refer to simply as "Baseline". Our reduced version achieves comparable or better performance with a similar or smaller number of parameters.
@@
On the Permuted MNIST benchmark (Table 3), SHIELD delivers the strongest adversarial robustness among all compared methods. It surpasses prior approaches under all attack settings, including AutoAttack, PGD, and FGSM, establishing itself as the most resilient model to adversarial perturbations in this setting. While a few baselines, particularly DGP and GPM, achieve higher clean accuracy, their robustness drops notably under stronger attacks. In contrast, SHIELD maintains both high robustness and solid performance on clean samples, without relying on overly large architectures or task-specific replay. This highlights the method’s ability to generalize under distribution shifts and adversarial pressure, a balance that remains a major challenge in continual learning.
Table 3: Comparison of AA after completing all tasks on the Permuted MNIST dataset. AA results for SHIELD are averaged over 2 seeds for AutoAttack and 5 seeds for all other evaluations.
@@
SHIELD $\mathbf{88.43\pm 0.35}$ $\mathbf{90.47\pm 1.36}$ $\mathbf{79.6\pm 1.34}$ 90.71±1.890.71$\pm$ 1.890.71 ± 1.8
Table 4: Comparison of AA after completing all tasks on the Rotated MNIST dataset. AA results for SHIELD are averaged over 2 seeds for AutoAttack and 5 seeds for all other evaluations.
@@
SHIELD $\mathbf{90.36\pm 0.1}$ $\mathbf{92.76\pm 0.22}$ $\mathbf{82.21\pm 0.43}$ 93.12±0.2293.12$\pm$ 0.2293.12 ± 0.22
On Rotated MNIST (Table 4), SHIELD sets a new state-of-the-art across all evaluated attacks. It achieves the highest Average Accuracy (AA) under AutoAttack (90.36%), PGD (92.76%), and FGSM (82.21%), outperforming strong baselines like DGP and GPM by a significant margin. Moreover, SHIELD not only maintains robust performance but also retains high AA (93.12%) on original samples, closely matching the top-performing baselines in that category.
Table 5: Comparison of AA after completing all tasks on the Split CIFAR-100 dataset. AA results for SHIELD are averaged over 2 seeds for AutoAttack and 5 seeds for all other evaluations.
@@
SHIELD $\mathbf{54.55\pm 0.42}$ $\mathbf{54.53\pm 0.48}$ 47.71±0.8647.71$\pm$ 0.8647.71 ± 0.86 59.36±1.1159.36$\pm$ 1.1159.36 ± 1.11
On the more complex Split CIFAR-100 benchmark (Table 5), SHIELD once again delivers the strongest performance under adversarial attacks. It outperforms all baselines by a wide margin in both AutoAttack and PGD evaluations, achieving 54.55% and 54.53% AA, respectively, substantially ahead of the next best method, DGP. While some baselines achieve slightly better AA on original samples, they fall short under adversarial threat models. In contrast, SHIELD consistently balances robustness and plasticity, confirming its capacity to learn durable representations without catastrophic forgetting, even in challenging continual learning settings.
@@
Appendix F Selected Hyperparameters
For the Split CIFAR-100 experiments, we used an AlexNet-based target network along with a hypernetwork comprising two hidden layers of 100 and 50 neurons, respectively. The embedding size was set to 512, with a batch size of 32 and a learning rate of 0.0010.0010.001. The hypernetwork used a regularization coefficient $\beta=0.01$, and training was performed with an $\ell_{\infty}$ perturbation of $\varepsilon=0.005$. The optimizer was Adam, and no data augmentation was used. Training was conducted for 200 epochs with ReLU activation in both the target and hypernetwork. We applied a learning rate scheduler: ReduceLROnPlateau with monitoring set to the maximum validation accuracy, reduction factor $\sqrt{0.1}$, patience of 5 epochs, minimum learning rate of $5\times 10^{-7}$, and no cooldown. Batch normalization was enabled throughout. When Interval MixUp is applied, we sample the mixing coefficient from a Beta distribution with parameter $\alpha=0.3$.
@@ Appendix F Selected Hyperparameters
In the Split miniImageNet experiments, we used a ResNet-18 target network and a hypernetwork with two hidden layers of 100 and 50 neurons, respectively. The embedding dimension was 128, with a batch size of 16 and a learning rate of 0.0010.0010.001. We trained using the Adam optimizer for 50 epochs, applying ReLU activations and enabling batch normalization throughout. The hypernetwork was regularized with $\beta=0.01$, and training included an $\ell_{\infty}$ perturbation of $\varepsilon=\frac{2}{255}$. No data augmentation was used. The learning rate scheduler matched that of the Split CIFAR-100 setup. When Interval MixUp was used, mixing coefficients were drawn from a Beta distribution with $\alpha=0.2$.
Model selection was based on the best validation loss. The optimizer across all experiments was Adam, and ReLU was used as the activation function throughout. In all cases, the same adversarial attack settings were used as in (Ru et al. 2024), except that the number of attack iterations for PGD was not specified in that work; we defaulted to 100 iterations for consistency across evaluations. For AutoAttack, we used the standard version with default parameters as suggested by the authors in (Croce and Hein 2020).
@@ Appendix G Training Details
Interval Bound Propagation Training Details
@@
$\varepsilon_{i}$ =$=\begin{cases}\frac{2i\cdot\varepsilon}{E}&\text{if }i\leq\lfloor\frac{E}{2}\rfloor,\\ \varepsilon&\text{otherwise}.\end{cases}$ (30)
Architecture Details
In the main part of the paper, we use the same base architecture across Permuted MNIST, Rotated MNIST, and Split-miniImageNet. For the Split CIFAR-100, we adopt a streamlined variant of AlexNet to meet the capacity and efficiency requirements of hypernetwork-based training. Specifically, we remove dropout, include biases, and replace standard batch normalization layers with interval-aware variants. The classifier comprises two fully connected layers with 100 units each. To improve numerical stability during IBP, we replace max pooling with average pooling. Additionally, a $\times 4$ 4 average pooling layer is applied after the final convolutional layer to reduce dimensionality.
This lightweight adaptation of AlexNet ensures that the target network remains compact, which is critical for training efficiency when its parameters are generated dynamically by a hypernetwork. Importantly, unlike (Ru et al. 2024), we do not allocate separate convolutional layers for each task; our model maintains a single shared architecture across tasks. Further architectural details and hyperparameter choices are discussed in Appendix Section G. A shared MLP hypernetwork is used across all datasets.
Appendix H Ablation Study
Alternative Epsilon Decay Rate Functions in Interval MixUp
@@ Appendix H Ablation Study
Figure 4: Comparison of different epsilon decay rates used in Interval MixUp.
Figure 5 presents the AA over time for various decay rate strategies across three continual learning benchmarks: Permuted MNIST, Split CIFAR-100, and Split miniImageNet. The top row shows AA, computed as the mean accuracy over all tasks seen so far, immediately after learning each task. The bottom row shows final AA, calculated as the average accuracy on all tasks after the final task is learned.
@@
Figure 5: Comparison of results for different epsilon decay rates across (a) Permuted MNIST, (b) Split CIFAR-100, and (c) Split miniImageNet. Top row: AA measured immediately after learning each task. Bottom row: AA evaluated on all tasks after completing the final task.
Table 6: Comparison of AA after completing all tasks on the Permuted MNIST dataset. AA results for different scheduling strategies (linear, quadratic, log, and cos) were calculated for a single seed.
@@
Table 7: Comparison of AA after completing all tasks on the Split CIFAR-100 dataset. AA results for different scheduling strategies (linear, quadratic, log, and cos) were calculated for a single seed.
@@
Table 8: Comparison of AA after completing all tasks on the Split miniImageNet dataset. AA results for different scheduling strategies (linear, quadratic, log, and cos) were calculated for a single seed.
@@
Impact of FGSM Perturbation Size on Average Accuracy
In Figure 6, we show the effect of FGSM attacks with varying $\varepsilon_{\text{attack}}$ values on the AA metric for our method and for HNET. The results indicate that HNET is substantially more vulnerable to adversarial perturbations, with a steep drop in AA on Split CIFAR-100 and Split miniImageNet, even under small $\varepsilon_{\text{attack}}$ values. For each subfigure in Figure 6, we evaluate our models using 10 FGSM attack strengths, with $\varepsilon_{\text{attack}}$ values evenly spaced from 0 up to twice the $\varepsilon$ (perturbation value used in the training process) used during training. This setup allows us to assess how well the models generalize robustness beyond the training-time perturbation level. The optimal $\varepsilon$ values used are those identified in Section F.
@@
Figure 6: AA of SHIELD and HNET under FGSM attacks with increasing perturbation sizes, evaluated on models trained on Permuted MNIST (Figure 6(a)), Split CIFAR-100 (Figure 6(b)), and Split miniImageNet (Figure 6(c)).
Impact of PGD Iterations on Average Accuracy
In Figure 7, we present the effect of PGD attacks with an increasing number of steps on the AA metric for SHIELD and HNET. Across all evaluated datasets, SHIELD consistently achieves higher initial accuracy when subjected to no attack or only a few PGD steps. Both SHIELD and HNET exhibit a noticeable drop in AA within the first 10 to 25 steps; however, their performance then stabilizes, demonstrating resilience to stronger iterative attacks.
For each subfigure in Figure 7, we evaluate both models across 10 PGD configurations, with the number of attack steps evenly spaced from 0 to 200. The remaining PGD hyperparameters (e.g., step size, $\varepsilon_{\text{attack}}$) are kept identical to those used in the main experiments. This extended evaluation allows us to assess model robustness across a broader attack range than previously reported. As before, we use the best-performing hyperparameters identified in Section F.

### v3

@@ 8 MixUp
9 Interval Neural Network Layers
@@
Interval Batch Normalization.
@@
More concretely, given pre-activation interval bounds $\underline{x},\overline{x}]$, we form the concatenated tensor:
@@
Interval Pooling Layers.
@@
and the resulting output interval is $\mathrm{MaxPool}_{\inf},\mathrm{MaxPool}_{\sup}]$.
@@
so that the final interval is $\mathrm{AvgPool}_{\inf},\mathrm{AvgPool}_{\sup}]$.
@@
10 Benchmark Datasets
Permuted MNIST introduces a different fixed pixel permutation for each task, applied to the original MNIST images. Rotated MNIST, on the other hand, creates tasks by rotating the digits by varying angles. These two datasets are widely used in the continual learning literature as synthetic benchmarks derived from MNIST [goodfellow2014explaining, liu2022continuallearningrecursivegradient]. Split CIFAR-100 [zenke2017continual] is constructed by randomly partitioning the 100 classes of the CIFAR-100 dataset into 10 mutually exclusive groups, with each group forming a separate classification task of 10 classes. Similarly, Split miniImageNet is generated by selecting a subset of the original ImageNet dataset [chaudhry2020continuallearninglowrankorthogonal] and dividing it into 20 distinct groups, each containing 5 unique classes. For both Split CIFAR-100 and Split miniImageNet, each class appears in only one group, ensuring no class overlap between tasks. This design allows each subset to be treated as an independent classification task, resulting in a sequence of tasks for continual learning. To match the setup of Split miniImageNet presented in [ru2024maintaining], we split Split miniImageNet as described; however, we use only the first ten tasks. Therefore, we have 10 tasks, each with 5 classes.
11 Discussion on Learnable Number of Parameters Used in SHIELD
@@
For the Permuted MNIST dataset, we reduce the number of learnable parameters to match the budget used in [ru2024maintaining] by employing a hypernetwork composed of a two-layer MLP with hidden dimensions 100 and 50, and an embedding size of 96. The target network is a two-layer MLP with 50 neurons per layer. For the Rotated MNIST dataset, we adopt the same architecture but reduce the embedding size to 24 to further minimize the parameter count. We use $\varepsilon=\frac{25}{255}$ for both Permuted MNIST and Rotated MNIST during training.
For the Split CIFAR-100 dataset, again, we align with the parameter budget from [ru2024maintaining] by using an embedding size of 512 and a hypernetwork consisting of a two-layer MLP with 200 and 50 hidden units, respectively. Additionally, we simplify the target AlexNet architecture by removing two fully connected layers, applying a $4\times 4$ average pooling operation after the final convolutional layer in place of the max pooling, and reducing the hidden dimension of the last linear layer to 100. For Split CIFAR-100, we use $\varepsilon=\frac{2}{255}$ during training.
The total number of learnable parameters used in each configuration is reported in Tab. 2. Furthermore, we adopt the same FGSM, PGD, and AutoAttack settings as described in the main paper.
Tabela 2: Comparison of learnable parameters (in millions) between our method and the baseline from [ru2024maintaining].
@@
On the Permuted MNIST benchmark (Tab. 3), SHIELD delivers the strongest adversarial robustness among all compared methods. It surpasses prior approaches under all attack settings, including AutoAttack, PGD, and FGSM, establishing itself as the most resilient model to adversarial perturbations in this setting. While a few baselines, particularly DGP and GPM, achieve higher clean accuracy, their robustness drops notably under stronger attacks. In contrast, SHIELD maintains both high robustness and solid performance on clean samples, without relying on overly large architectures or task-specific replay. This highlights the method’s ability to generalize under distribution shifts and adversarial pressure, a balance that remains a major challenge in continual learning.
Tabela 3: Comparison of AA after completing all tasks on the Permuted MNIST dataset. AA results for SHIELD are averaged over 2 seeds for AutoAttack and 5 seeds for all other evaluations.
@@
SHIELD 88.43 $\pm$ 0.35 90.47 $\pm$ 1.36 79.6 $\pm$ 1.34 90.71 $\pm$ 1.8
Tabela 4: Comparison of AA after completing all tasks on the Rotated MNIST dataset. AA results for SHIELD are averaged over 2 seeds for AutoAttack and 5 seeds for all other evaluations.
@@
SHIELD 90.36 $\pm$ 0.1 92.76 $\pm$ 0.22 82.21 $\pm$ 0.43 $93.12\pm 0.22$
On Rotated MNIST (Tab. 4), SHIELD sets a new state-of-the-art across all evaluated attacks. It achieves the highest Average Accuracy (AA) under AutoAttack (90.36%), PGD (92.76%), and FGSM (82.21%), outperforming strong baselines like DGP and GPM by a significant margin. Moreover, SHIELD not only maintains robust performance but also retains high AA (93.12%) on original samples, closely matching the top-performing baselines in that category.
Tabela 5: Comparison of AA after completing all tasks on the Split CIFAR-100 dataset. AA results for SHIELD are averaged over 2 seeds for AutoAttack and 5 seeds for all other evaluations.
@@
SHIELD 54.55 $\pm$ 0.42 54.53 $\pm$ 0.48 47.71 $\pm$ 0.86 59.36 $\pm$ 1.11
On the more complex Split CIFAR-100 benchmark (Tab. 5), SHIELD once again delivers the strongest performance under adversarial attacks. It outperforms all baselines by a wide margin in both AutoAttack and PGD evaluations, achieving 54.55% and 54.53% AA, respectively, substantially ahead of the next best method, DGP. While some baselines achieve slightly better AA on original samples, they fall short under adversarial threat models. In contrast, SHIELD consistently balances robustness and plasticity, confirming its capacity to learn durable representations without catastrophic forgetting, even in challenging continual learning settings.
@@
12 Selected Hyperparameters
For the Split CIFAR-100 experiments, we used an AlexNet-based target network along with a hypernetwork comprising two hidden layers of 100 and 50 neurons, respectively. The embedding size was set to 512, with a batch size of 32 and a learning rate of $0.001$. The hypernetwork used a regularization coefficient $\beta=0.01$, and training was performed with an $\ell_{\infty}$ perturbation of $\varepsilon=0.005$. The optimizer was Adam, and no data augmentation was used. Training was conducted for 200 epochs with ReLU activation in both the target and hypernetwork. We applied a learning rate scheduler: ReduceLROnPlateau with monitoring set to the maximum validation accuracy, reduction factor $\sqrt{0.1}$, patience of 5 epochs, minimum learning rate of $5\times 10^{-7}$, and no cooldown. Batch normalization was enabled throughout. When Interval MixUp is applied, we sample the mixing coefficient from a Beta distribution with parameter $\alpha=0.3$.
@@ 12 Selected Hyperparameters
In the Split miniImageNet experiments, we used a ResNet-18 target network and a hypernetwork with two hidden layers of 100 and 50 neurons, respectively. The embedding dimension was 128, with a batch size of 16 and a learning rate of $0.001$. We trained using the Adam optimizer for 50 epochs, applying ReLU activations and enabling batch normalization throughout. The hypernetwork was regularized with $\beta=0.01$, and training included an $\ell_{\infty}$ perturbation of $\varepsilon=\frac{2}{255}$. No data augmentation was used. The learning rate scheduler matched that of the Split CIFAR-100 setup. When Interval MixUp was used, mixing coefficients were drawn from a Beta distribution with $\alpha=0.2$.
Model selection was based on the best validation loss. The optimizer across all experiments was Adam, and ReLU was used as the activation function throughout. In all cases, the same adversarial attack settings were used as in [ru2024maintaining], except that the number of attack iterations for PGD was not specified in that work; we defaulted to 100 iterations for consistency across evaluations. For AutoAttack, we used the standard version with default parameters as suggested by the authors in [croce2020reliableevaluationadversarialrobustness].
@@ 13 Training Details
Interval Bound Propagation Training Details.
@@
$\varepsilon_{i}$ $=\begin{cases}\frac{2i\cdot\varepsilon}{E}&\text{if }i\leq\lfloor\frac{E}{2}\rfloor,\\ \varepsilon&\text{otherwise}.\end{cases}$ (31)
Architecture Details.
In the main part of the paper, we use the same base architecture across Permuted MNIST, Rotated MNIST, and Split-miniImageNet. For the Split CIFAR-100, we adopt a streamlined variant of AlexNet to meet the capacity and efficiency requirements of hypernetwork-based training. Specifically, we remove dropout, include biases, and replace standard batch normalization layers with interval-aware variants. The classifier comprises two fully connected layers with 100 units each. To improve numerical stability during IBP, we replace max pooling with average pooling. Additionally, a $4\times 4$ average pooling layer is applied after the final convolutional layer to reduce dimensionality.
This lightweight adaptation of AlexNet ensures that the target network remains compact, which is critical for training efficiency when its parameters are generated dynamically by a hypernetwork. Importantly, unlike [ru2024maintaining], we do not allocate separate convolutional layers for each task; our model maintains a single shared architecture across tasks. Further architectural details and hyperparameter choices are discussed in Appendix Section 13. A shared MLP hypernetwork is used across all datasets.
14 Ablation Study and Additional Experiments
14.1 Alternative Epsilon Decay Rate Functions in Interval MixUp
@@
Rysunek 4: Comparison of different epsilon decay rates used in Interval MixUp.
Fig. 5 presents the AA over time for various decay rate strategies across three continual learning benchmarks: Permuted MNIST, Split CIFAR-100, and Split miniImageNet. The top row shows AA, computed as the mean accuracy over all tasks seen so far, immediately after learning each task. The bottom row shows final AA, calculated as the average accuracy on all tasks after the final task is learned.
@@
Rysunek 5: Comparison of results for different epsilon decay rates across (a) Permuted MNIST, (b) Split CIFAR-100, and (c) Split miniImageNet. Top row: AA measured immediately after learning each task. Bottom row: AA evaluated on all tasks after completing the final task.
Tabela 6: Comparison of AA after completing all tasks on the Permuted MNIST dataset. AA results for different scheduling strategies (linear, quadratic, log, and cos) were calculated for a single seed.
@@
Tabela 7: Comparison of AA after completing all tasks on the Split CIFAR-100 dataset. AA results for different scheduling strategies (linear, quadratic, log, and cos) were calculated for a single seed.
@@
Tabela 8: Comparison of AA after completing all tasks on the Split miniImageNet dataset. AA results for different scheduling strategies (linear, quadratic, log, and cos) were calculated for a single seed.
@@
14.2 Impact of FGSM Perturbation Size on Average Accuracy
In Fig. 6, we show the effect of FGSM attacks with varying $\varepsilon_{\text{attack}}$ values on the AA metric for our method and for HNET. The results indicate that HNET is substantially more vulnerable to adversarial perturbations, with a steep drop in AA on Split CIFAR-100 and Split miniImageNet, even under small $\varepsilon_{\text{attack}}$ values. For each subfigure in Fig. 6, we evaluate our models using 10 FGSM attack strengths, with $\varepsilon_{\text{attack}}$ values evenly spaced from 0 up to twice the $\varepsilon$ (perturbation value used in the training process) used during training. This setup allows us to assess how well the models generalize robustness beyond the training-time perturbation level. The optimal $\varepsilon$ values used are those identified in Section 12.
@@ 14.2 Impact of FGSM Perturbation Size on Average Accuracy
Rysunek 6: AA of SHIELD and HNET under FGSM attacks with increasing perturbation sizes, evaluated on models trained on Permuted MNIST (Fig. 6(a)), Split CIFAR-100 (Fig. 6(b)), and Split miniImageNet (Fig. 6(c)).
14.3 Impact of PGD Iterations on Average Accuracy
In Fig. 7, we present the effect of PGD attacks with an increasing number of steps on the AA metric for SHIELD and HNET. Across all evaluated datasets, SHIELD consistently achieves higher initial accuracy when subjected to no attack or only a few PGD steps. Both SHIELD and HNET exhibit a noticeable drop in AA within the first 10 to 25 steps; however, their performance then stabilizes, demonstrating resilience to stronger iterative attacks.
For each subfigure in Fig. 7, we evaluate both models across 10 PGD configurations, with the number of attack steps evenly spaced from 0 to 200. The remaining PGD hyperparameters (e.g., step size, $\varepsilon_{\text{attack}}$) are kept identical to those used in the main experiments. This extended evaluation allows us to assess model robustness across a broader attack range than previously reported. As before, we use the best-performing hyperparameters identified in Section 12.